OpenStudy (narissa):

Please help 2x+6-6x=30 A.-6 B.-3 C.3 D.6

OpenStudy (vera_ewing):

Ok, so first you need to isolate the variable.

OpenStudy (vera_ewing):

In this case, the variable is \[x\]

OpenStudy (vera_ewing):

So, you would get 2x+6-6x=30 2x-6x = 30-6 -4x = 24

OpenStudy (narissa):

so i subtract -4x from 2x-6x right?

OpenStudy (vera_ewing):

Now, solve: -4x = 24

OpenStudy (vera_ewing):